### v1

# Size-consistent implementation of Hamiltonian simulation-based quantum-selected configuration interaction method for the supramolecular approach

Abstract
The quantum-selected configuration interaction (QSCI) method is a promising approach for large-scale quantum chemical calculations on currently available quantum hardware. However, its naive implementation lacks size consistency, which is essential for accurate intermolecular interaction energy calculations using the supramolecular approach. Here, we present a size-consistent implementation of QSCI by sampling Slater determinants for the dimer in the localized molecular orbital basis, constructing the subspaces for the monomers and dimer, and augmenting the dimer subspace with additional determinants required for size consistency. Implemented within the Hamiltonian simulation-based QSCI (HSB-QSCI) framework, our method numerically satisfies size consistency for 4H/8H clusters, the FH dimer, and the FH–H2O system. Application to intermolecular interaction energy calculations of hydrogen-bonded FH dimer and FH–H2O demonstrates that our approach reproduces complete active space-configuration interaction (CAS-CI) values with errors below 0.04 kcal mol-1.
I Introduction
Quantum computation holds the potential to revolutionize our society. Among its diverse applications, sophisticated quantum chemical calculations are considered one of the most promising. Accurate quantum chemical calculations can accelerate drug and materials discovery, elucidate reaction mechanisms in enzymes and biomolecules, and so on. Quantum phase estimation (QPE) [1, 2, 3] and variational quantum eigensolver (VQE) [4] are two major quantum computational approaches, targeting the fault-tolerant quantum computing (FTQC) and noisy intermediate-scale quantum (NISQ) eras, respectively, and both methods have been extensively investigated from theoretical and hardware demonstration perspectives. However, QPE typically requires quantum circuit that are prohibitively deep for even the most advanced quantum hardware available today, rendering its application to industrially relevant molecules extremely challenging until FTQC becomes a reality. On the other hand, VQE can be executed on NISQ devices, but it faces several significant challenges, such as substantial measurement overhead [5], the presence of barren plateaus [6], and so on.
The quantum-selected configuration interaction (QSCI) method was proposed to perform large-scale quantum chemical calculations using currently available quantum hardware [7]. In QSCI, a quantum computer is used to sample Slater determinants that make significant contributions to the wave function of the target electronic state. The information from these sampled Slater determinants is then transferred to a classical computer, where the subspace Hamiltonian is constructed and diagonalized to obtain the energy and wave function. The key step in QSCI is the preparation of a quantum state suitable for sampling important Slater determinants. Initially, quantum state preparation methods based on VQE and adiabatic state preparation were proposed for this purpose [7]. In 2025, Robledo-Moreno and coworkers introduced a method for quantum state preparation based on the local unitary cluster Jastrow (LUCJ) ansatz [8] with classically computed coupled cluster singles and doubles (CCSD) excitation amplitudes. They demonstrated QSCI calculations using up to 77 qubits on superconducting quantum processors, aided by an error mitigation technique called self-consistent configuration recovery [9]. Recently, we proposed a quantum state preparation method for QSCI based on Hamiltonian simulation (HSB-QSCI) [10], in which important Slater determinants are sampled from quantum states generated by real-time evolution of an approximate wave function. HSB-QSCI offers several advantages over the original QSCI: (1) a simple wave function such as a Hartree–Fock (HF) wave function can be used as the initial state for time evolution, (2) VQE variational optimization is not required, (3) different samples can be obtained by varying the evolution time, (4) long-time evolution is challenging to simulate on classical computers, and (5) higher-order excited configurations can be sampled even with shorter evolution times. We have also reported proof-of-principle demonstrations of HSB-QSCI for one-dimensional carbon chain molecules (carbynes), utilizing up to 36 qubits of an IBM superconducting quantum processor, with the aid of matrix product operator (MPO)-based classical optimization of the quantum circuit for the time evolution operator [11]. Shortly after our publication, methods based on the same concept were reported by other groups [12, 13, 14]. Notably, Weaving and coworkers reported that time evolution-based state preparation for QSCI yields a more compact wave function expansion than conventional selected CI in SiH4 system [15].
Although QSCI has been actively investigated from both theoretical and experimental perspectives [16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32], it suffers from a serious drawback: QSCI is not size consistent unless special care is taken. In its naive implementation, the energy of a dimer composed of non-interacting monomers is generally not equal to the sum of the energies of the individual monomers. In fact, as discussed in the Supporting Information, the number of quantum circuit executions must increase exponentially with system size even approximately satisfy the size consistency condition. Consequently, the straightforward application of QSCI to the calculation of intermolecular interaction energies using the supramolecular approach is fundamentally impractical. In the supramolecular approach, the intermolecular interaction energy $E_{\text{int}}$ is calculated using the following equation:
$E_{\text{int}}=E_{\text{AB}}-E_{\text{A}}-E_{\text{B}}$ (1)
Here, $E_{\text{A}}$, $E_{\text{B}}$, and $E_{\text{AB}}$ represent the energies of monomer A, monomer B, and the interacting dimer (A + B), respectively. To address the challenges arising from the lack of size consistency in QSCI, the intermolecular interaction energy has instead been calculated using the following equation,
$E_{\text{int}}=E_{\text{AB}}-E_{\text{A}\cdots\text{B}}$ (2)
Here, $E_{\text{A}\cdots\text{B}}$ denotes the energy of the dimer composed of monomers A and B at a sufficiently large inter-monomer distance. Hereafter, we refer to the approach based on equation (2) as the “dimer approach”. In 2025, Kaliakin and coworkers reported QSCI-based calculations of intermolecular interaction energies for water and methane dimers using the dimer approach [19]. However, it is evident that the computational cost is lower for the supramolecular approach than for the dimer approach, since the size of the subspace Hamiltonian to be diagonalized is smaller for monomers than for the dimer. It is also important to note that if Slater determinants are sampled independently for different geometries, the electronic configurations included in the subspace Hamiltonian may differ between geometries, potentially resulting in discontinuities in the potential energy curve. In this work, we propose a size-consistent implementation of HSB-QSCI (sc-HSB-QSCI) to overcome these challenges.
II Theory
In conventional HSB-QSCI [10], Slater determinants used for subspace Hamiltonian diagonalization are sampled from quantum states generated by real-time evolution of an approximate wave function. To expand the subspace, measurement outcomes obtained at different evolution times, $k\Delta t$ for $k\leq K$, are combined. For example, HSB-QSCI with $K=3$ indicates that the subspace Hamiltonian is constructed using the measurement results from quantum circuit executions at $k$ = 1, 2, and 3.
Figure 1: Schematic illustration of subspace generation in sc-HSB-QSCI. $|\Phi_{\text{Di}}\rangle$ represents an approximate wave function for the dimer, and $U$ is a time evolution operator $U=e^{-iHt}$.
A schematic illustration of subspace construction in sc-HSB-QSCI is shown in Figure 1. Note that the following subspace construction approach is not limited to HSB-QSCI and can be universally applied to any selected CI framework. The strategy of sc-HSB-QSCI is as follows: (1) Sample Slater determinants for the dimer in the localized molecular orbital basis. (2) For each sampled bit string, check the electronic configuration of the dimer. If the configuration involves only intra-monomer excitations, split the bit string into those corresponding to each monomer and add them to the respective monomer subspace in the localized molecular orbital basis. If inter-monomer charge-transfer excitations are present, add the bit string to the dimer subspace without splitting. (3) Perform subspace Hamiltonian diagonalization for the monomers using the Slater determinants generated in the previous step. (4) The dimer subspace Hamiltonian is constructed by including all possible combinations of the electronic configurations from monomer 1 and monomer 2 subspaces, as well as the Slater determinants corresponding to charge-transfer excitations. With this approach, electronic configurations required for size consistency but not sampled are systematically added, thereby ensuring that size consistency is automatically satisfied.
III Computational conditions
As a proof-of-principle demonstration of sc-HSB-QSCI, we performed numerical simulations on a classical computer. The target systems included a dimer of square tetrahydrogen (4H) clusters and hydrogen-bonded systems (FH dimer and FH–H2O). The H–H bond length in the 4H cluster was set to 1.0583 Å (= 2.0 Bohr), and the 8H cluster was constructed by placing two 4H clusters in a cuboid arrangement with an inter-monomer distance of 100 Å. Note that this system has previously been used to investigate size consistency in VQE with a unitary coupled cluster singles and doubles (UCCSD) ansatz [33], and in QPE [34]. The localized molecular orbitals of the 8H cluster were manually generated by taking appropriate linear combinations of RHF/STO-3G [35] canonical orbitals. For sc-HSB-QSCI, all molecular orbitals were included in the Hamiltonian simulation, and the interaction energies were compared with those from full-configuration interaction (full-CI) calculations.
Figure 2: PBE0/aug-$c$-pVDZ optimized geometry of the FH dimer and FH–H2O. Red dotted lines represent hydrogen bonds, and values in red denote intermolecular distances in units of Å.
For the hydrogen-bonded systems, geometry optimizations of the dimers were performed at the PBE0 [36]/aug-cc-pVDZ [37, 38] level of theory. No imaginary vibrational frequencies were obtained at the optimized structure. The optimized geometries are shown in Figure 2, and the corresponding Cartesian coordinates are provided in Tables S1 and S2 in the Supporting Information. For monomer calculations, we used the optimized geometry for the dimer without further performing geometry optimization of the monomers. To assess size consistency, we also generated geometries for the dimers without inter-monomer interaction by increasing the inter-monomer distance to 100 Å. Localized molecular orbitals for both dimers and monomers were generated using the Pipek–Mezey method [39] at the RHF/aug-cc-pVDZ level. In the sc-HSB-QSCI simulations, only valence orbitals were considered, with active space sizes of (16e, 10o) for the FH dimer, (8e, 5o) for the FH monomer, (16e, 11o) for the FH–H2O, and (8e, 6o) for the H2O. Here, (NNe, MMo) indicates an active space consisting of $N$ electrons and $M$ molecular orbitals. The active orbitals for the monomers and dimers of the FH dimer and FH–H2O are shown in Figure S2 in Supporting Information. All DFT and RHF calculations were performed using the GAMESS-US program package [40].
Numerical simulations for sc-HSB-QSCI were performed using our in-house Python3 code, developed with the OpenFermion [41], Cirq [42], and PyCI [43] libraries. Details of the HSB-QSCI implementation can be found in Reference [10]. The Jordan–Wigner transformation [44] was used for fermion–qubit encoding. The RHF wave function served as the initial state for time evolution, and the quantum circuit for the time evolution operator $U=e^{-iH\Delta t}$ with $\Delta t=1$ was constructed using the second-order Trotter–Suzuki decomposition with a single Trotter step. Unless otherwise noted, the number of shots for quantum circuit execution was set to $1\times 10^{4}$, and $K$ is varied from 1 to 5 and 10. In sc-HSB-QSCI, symmetry completion [10] was performed prior to constructing the subspace Hamiltonian matrix. In this step, any open-shell Slater determinants required to span the eigenfunctions of the electron spin ${\bf S}^{2}$ operator were added if missing.
For the hydrogen-bonded FH dimer and FH–H2O systems, we also calculated intermolecular interaction energies using HCI in conjunction with the propose approach. The HCI calculations were performed using the PyCI [43] library.
IV Results and Discussion
IV.1 Size consistency of sc-HSB-QSCI
Table 1: Sc-HSB-QSCI, RHF, and full-CI energies for the 8H (dimer) and 4H (monomer) clusters, in units of Hartree. The numbers in parenthesis indicate the number of Slater determinants included in the subspace Hamiltonian diagonalization.
$K$ EE(dimer) EE(monomer 1) EE(monomer 2)
1 -3.652202 (98) -1.816720 (8) -1.835483 (8)
2 -3.862635 (124) -1.933870 (9) -1.928764 (10)
3 -3.862635 (124) -1.933870 (9) -1.928764 (10)
4 -3.862664 (134) -1.933900 (10) -1.928764 (10)
5 -3.862664 (134) -1.933900 (10) -1.928764 (10)
RHF -3.553541 (1) -1.776770 (1) -1.776770 (1)
full-CI -3.878863 (178) -1.939432 (12) -1.939432 (12)
Here, we numerically demonstrate that the sc-HSB-QSCI method satisfies the size consistency condition, using the 4H cluster as the monomer. The square 4H cluster is a well-known strongly correlated system [45], with a ground-state wave function exhibiting significant multi-configurational character. Using the STO-3G basis set [35], the full-CI dimensions are 12 and 178 for the monomer (4H cluster in the $D_{2h}$ point group) and the dimer (8H cluster in the $C_{2v}$ point group), respectively. Owing to the small system size, sc-HSB-QSCI calculations were performed with 10 shots for each $k$. The results are summarized in Table 1. The computed $E_{\text{int}}$ values are less than $1\times 10^{-7}$ Hartree for all $K$. Tne small deviations of $E_{\text{int}}$ from zero arize from rounding errors in the one- and two-electron integrals. These results clearly demonstrate the size consistent nature of sc-HSB-QSCI.
Table 2: Intermolecular interaction energies for the FH dimer and FH–H2O at an inter-monomer distance of 100 Å, calculated using sc-HSB-QSCI and org-HSB-QSCI, in units of kcal mol-1.
Molecule $K$ sc-HSB-QSCI org-HSB-QSCI
FH dimer 1 0.000 0.016
2 0.000 0.016
3 0.000 0.016
4 0.000 0.013
5 0.000 0.007
FH–H2O 1 0.000 0.095
2 0.000 0.046
3 0.000 0.040
4 0.000 0.040
5 0.000 0.031
Next, we numerically investigated size consistency in the FH dimer and FH–H2O systems using both sc-HSB-QSCI and the original form of HSB-QSCI (org-HSB-QSCI). As discussed above, org-HSB-QSCI is not size consistent, and the interaction energy can be nonzero even when the intermolecular distance is sufficiently large. In this study, we used geometries with an intermolecular distance of 100 Å. The calculated interaction energies are summarized in Table 2, and total energies and the number of Slater determinants in the subspace are provided in Tables S3 and S4 in the Supporting Information. Sc-HSB-QSCI yielded an interaction energy $E_{\text{int}}$ = 0.000 kcal mol-1 for all $K$, whereas org-HSB-QSCI gave nonzero interaction energies. The error in $E_{\text{int}}$ is slightly larger for FH–H2O than for the FH dimer, suggesting that the violation of size consistency can increase as the number of active orbitals grows. In all systems under study, the number of sampled charge-transfer excited Slater determinants was zero.
IV.2 Intermolecular interaction energies of hydrogen-bonded FH dimer and FH–H2O
Lastly, we examine the intermolecular interaction energy calculations of the hydrogen-bonded FH dimer and FH–H2O. Hydrogen bonding is one of the most important intermolecular interactions, particularly in biomolecules. Here, we compare interaction energies calculated using the supramolecular approach with sc-HSB-QSCI and org-HSB-QSCI, as well as the dimer approach with org-HSB-QSCI. The numerical simulation results, together with RHF and CAS-CI interaction energies, are summarized in Table 3, and the total energies are listed in Tables S5 and S6 in the Supporting Information. Sc-HSB-QSCI yields larger interaction energies in absolute value and smaller errors relative to the CAS-CI values compared to org-HSB-QSCI. Org-HSB-QSCI using the dimer approach gave interaction energies comparable to those obtained with sc-HSB-QSCI, although its computational cost is higher than that of the supramolecular approach. At $K=10$, sc-HSB-QSCI gave interaction energies with deviations of 0.012 and 0.038 kcal mol-1 from the CAS-CI values for the FH dimer and FH–H2O, respectively. The number of Slater determinants included in the subspace Hamiltonian diagonalization for the dimer is 458 and 4150 for the FH dimer and FH–H2O, respectively, which should be compared to 2025 and 27225, respectively, at the CAS-CI level of theory.
Table S7 in the Supporting Information presents the classification of bit strings sampled by the quantum computer into those corresponding to intra-monomer excited determinants and those representing inter-monomer charge-transfer excited determinants. Whether intra-monomer excitation or inter-monomer charge-transfer excitation is sampled more frequently depends on the duration of the time evolution. As $k$ increases, the probability of sampling charge-transfer excitation becomes higher. This indicates that longer time evolution is important for accurately obtaining intermolecular interaction energies using sc-HSB-QSCI.
Table 3: Intermolecular interaction energies of the hydrogen-bonded FH dimer and FH–H2O, calculated using the supramolecular approach with sc-HSB-QSCI and org-HSB-QSCI, as well as the dimer approach with org-HSB-QSCI, in units of kcal mol-1.
System $K$ Supramolecular approach Dimer approach RHF CAS-CI
sc-HSB-QSCI org-HSB-QSCI org-HSB-QSCI
FH dimer 1 4.030-4.030 4.002-4.002 4.019-4.019 3.971-3.971 4.071-4.071
2 4.037-4.037 4.011-4.011 4.027-4.027
3 4.043-4.043 4.020-4.020 4.036-4.036
4 4.055-4.055 4.040-4.040 4.054-4.054
5 4.056-4.056 4.041-4.041 4.048-4.048
10 4.059-4.059 4.049-4.049 4.053-4.053
FH–H2O 1 8.450-8.450 8.140-8.140 8.234-8.234 7.902-7.902 8.786-8.786
2 8.497-8.497 8.402-8.402 8.447-8.447
3 8.687-8.687 8.600-8.600 8.640-8.640
4 8.695-8.695 8.617-8.617 8.658-8.658
5 8.717-8.717 8.674-8.674 8.705-8.705
10 8.748-8.748 8.717-8.717 8.741-8.741
Since our size-consistent implementation is applicable to other selected CI frameworks, we applied our proposed scheme to heat-bath CI (HCI) [46]. Details of the numerical simulations are provided in subsection C of the Supporting Information. The size consistent HCI (sc-HCI) yielded intermolecular interaction energies slightly closer to the CAS-CI values than those obtained with sc-HSB-QSCI, indicating that sc-HCI achieves a more compact wave function expansion for a given level of accuracy. As discussed in our previous work [10], HSB-QSCI may sample Slater determinants that are important for excited states rather than the ground state. Therefore, screening out unimportant Slater determinants from the subspace is essential for obtaining a more compact wave function expansion in sc-HSB-QSCI, which remains one of the most significant challenges in QSCI. Nevertheless, the approach proposed in this study significantly improves the applicability of selected CI methods to intermolecular interaction energy calculations.
V Conclusion
In summary, we have developed a size-consistent implementation of Hamiltonian simulation-based quantum-selected configuration interaction (sc-HSB-QSCI) for intermolecular interaction energy calculations based on the supramolecular approach. The size consistency of the proposed method was confirmed through numerical simulations of the 8H cluster, FH dimer, and FH–H2O systems with an inter-monomer distance of 100 Å, all yielding zero interaction energies. Interaction energies of intermolecular hydrogen bonds in the FH dimer and FH–H2O were also calculated with sc-HSB-QSCI, and the results agreed with the CAS-CI interaction energies in the same active space with errors of only 0.012 and 0.038 kcal mol-1 for the FH dimer and FH–H2O, respectively. During the sampling procedure, intra-monomer excited configurations are predominantly sampled in the early stage of real-time evolution, while the probability of sampling charge-transfer excited configurations increases at later stages. This trend indicates that extended real-time evolution is essential for accurately computing intermolecular interaction energies with sc-HSB-QSCI.
Apart from the supramolecular approach, size consistency is also crucial in other contexts, such as the fragment molecular orbital method [47, 48] and the ONIOM (our own N-layer integrated molecular orbitals and molecular mechanics) method [49, 50]. Applying the sc-HSB-QSCI method to these problems remains an important avenue for future research.
VI Acknowledgments
K.S. thanks Dr. Yuhei Tachi and Dr. Hiroyuki Tezuka for useful discussions.
D. Localized molecular orbitals and sc-HSB-QSCI results Refer to caption
Figure S2: Localized molecular orbitals of FH dimer and FH–H2O used as active space for HSB-QSCI. Orbitals with yellow highlights are occupied ones, and those without highlight are unoccupied.
Table S3: Sc-HSB-QSCI, org-HSB-QSCI, RHF, and full-CI energies of the FH dimer at an intermolecular distance of 100 Å, in units of Hartree. The numbers in parenthesis indicate the number of Slater determinants in the subspace.
$K$ sc-HSB-QSCI org-HSB-QSCI
EE(dimer) EE(monomer 1) EE(monomer 2) EE(dimer) EE(monomer 1) EE(monomer 2)
1 200.097212-200.097212 (139) 100.048655-100.048655 (11) 100.048558-100.048558 (11) 200.097187-200.097187 (39) 100.048655-100.048655 (11) 100.048558-100.048558 (11)
2 200.097212-200.097212 (139) 100.048655-100.048655 (11) 100.048558-100.048558 (11) 200.097187-200.097187 (48) 100.048655-100.048655 (11) 100.048558-100.048558 (11)
3 200.097212-200.097212 (139) 100.048655-100.048655 (11) 100.048558-100.048558 (11) 200.097187-200.097187 (48) 100.048655-100.048655 (11) 100.048558-100.048558 (11)
4 200.097212-200.097212 (139) 100.048655-100.048655 (11) 100.048558-100.048558 (11) 200.097191-200.097191 (52) 100.048655-100.048655 (11) 100.048558-100.048558 (11)
5 200.097212-200.097212 (139) 100.048655-100.048655 (11) 100.048558-100.048558 (11) 200.097202-200.097202 (62) 100.048655-100.048655 (11) 100.048558-100.048558 (11)
RHF 200.065881-200.065881 (1) 100.033096-100.033096 (1) 100.032785-100.032785 (1) 200.065881-200.065881 (1) 100.033096-100.033096 (1) 100.032785-100.032785 (1)
full-CI 200.097212-200.097212 (2025) 100.048655-100.048655 (25) 100.048558-100.048558 (25) 200.097212-200.097212 (2025) 100.048655-100.048655 (25) 100.048558-100.048558 (25)
Table S4: Sc-HSB-QSCI, org-HSB-QSCI, RHF, and full-CI energies of the FH–H2O at an intermolecular distance of 100 Å, in units of Hartree. The numbers in parenthesis indicate the number of Slater determinants in the subspace.
$K$ sc-HSB-QSCI org-HSB-QSCI
EE(dimer) EE(monomer 1) EE(monomer 2) EE(dimer) EE(monomer 1) EE(monomer 2)
1 176.128580-176.128580 (1327) 76.080556-76.080556 (95) 100.048023-100.048023 (11) 176.128507-176.128507 (198) 76.080634-76.080634 (86) 100.048023-100.048023 (11)
2 176.128687-176.128687 (1394) 76.080663-76.080663 (100) 100.048023-100.048023 (11) 176.128616-176.128616 (213) 76.080665-76.080665 (99) 100.048023-100.048023 (11)
3 176.128697-176.128697 (1433) 76.080674-76.080674 (103) 100.048023-100.048023 (11) 176.128626-176.128626 (218) 76.080667-76.080667 (107) 100.048023-100.048023 (11)
4 176.128703-176.128703 (1545) 76.080680-76.080680 (111) 100.048023-100.048023 (11) 176.128641-176.128641 (246) 76.080683-76.080683 (119) 100.048023-100.048023 (11)
5 176.128710-176.128710 (1629) 76.080686-76.080686 (117) 100.048023-100.048023 (11) 176.128661-176.128661 (301) 76.080687-76.080687 (121) 100.048023-100.048023 (11)
RHF 176.072736-176.072736 (1) 76.041090-76.041090 (1) 100.031644-100.031644 (1) 176.072736-176.072736 (1) 76.041090-76.041090 (1) 100.031644-100.031644 (1)
full-CI 176.128718-176.128718 (27225) 76.080692-76.080692 (225) 100.048024-100.048024 (25) 176.128718-176.128718 (27225) 76.080692-76.080692 (225) 100.048024-100.048024 (25)
Table S5: Sc-HSB-QSCI, org-HSB-QSCI, RHF, and full-CI energies of the hydrogen-bonded FH dimer, in units of Hartree. The numbers in parenthesis indicate the number of Slater determinants in the subspace.
$K$ sc-HSB-QSCI org-HSB-QSCI
EE(dimer) EE(monomer 1) EE(monomer 2) EE(dimer)
1 200.103634-200.103634 (303) 100.048655-100.048655 (15) 100.048558-100.048558 (13) 200.103590-200.103590 (134)
2 200.103645-200.103645 (311) 100.048655-100.048655 (15) 100.048558-100.048558 (13) 200.103605-200.103605 (156)
3 200.103655-200.103655 (318) 100.048655-100.048655 (15) 100.048558-100.048558 (13) 200.103618-200.103618 (178)
4 200.103674-200.103674 (410) 100.048655-100.048655 (17) 100.048558-100.048558 (15) 200.103651-200.103651 (204)
5 200.103676-200.103676 (416) 100.048655-100.048655 (17) 100.048558-100.048558 (15) 200.103653-200.103653 (216)
10 200.103681-200.103681 (458) 100.048655-100.048655 (17) 100.048558-100.048558 (15) 200.103665-200.103665 (280)
RHF 200.072209-200.072209 (1) 100.033096-100.033096 (1) 100.032785-100.032785 (1) 200.072209-200.072209 (1)
full-CI 200.103699-200.103699 (2025) 100.048655-100.048655 (25) 100.048558-100.048558 (25) 200.103699-200.103699 (2025)
Table S6: Sc-HSB-QSCI, org-HSB-QSCI, RHF, and full-CI energies of the hydrogen-bonded FH–H2O, in units of Hartree. The numbers in parenthesis indicate the number of Slater determinants in the subspace.
$K$ sc-HSB-QSCI org-HSB-QSCI
EE(dimer) EE(monomer 1) EE(monomer 2) EE(dimer)
1 176.141994-176.141994 (1608) 76.080505-76.080505 (97) 100.048023-100.048023 (11) 176.141629-176.141629 (514)
2 176.142184-176.142184 (1911) 76.080621-76.080621 (112) 100.048023-100.048023 (11) 176.142078-176.142078 (685)
3 176.142490-176.142490 (2293) 76.080624-76.080624 (132) 100.048023-100.048023 (11) 176.142395-176.142395 (861)
4 176.142513-176.142513 (2907) 76.080634-76.080634 (140) 100.048023-100.048023 (13) 176.142439-176.142439 (985)
5 176.142601-176.142601 (3239) 76.080686-76.080686 (151) 100.048023-100.048023 (13) 176.142533-176.142533 (1197)
10 176.142652-176.142652 (4150) 76.080688-76.080688 (184) 100.048023-100.048023 (13) 176.142607-176.142607 (1789)
RHF 176.085325-176.085325 (1) 76.041090-76.041090 (1) 100.031644-100.031644 (1) 176.085325-176.085325 (1)
full-CI 176.142717-176.142717 (27225) 76.080693-76.080693 (225) 100.048023-100.048023 (25) 176.142717-176.142717 (27225)
Table S7: Number of sampled bit strings corresponding to intra-monomer excited and inter-monomer charge-transfer excited Slater determinants.
Molecule $K$ Intra-monomer Charge transfer
FH dimer 1 40 37
2 46 51
3 54 60
4 66 74
5 68 77
10 87 102
FH–H2O 1 117 84
2 151 113
3 184 160
4 210 199
5 235 241
10 328 409